OpenStudy (isabel☺):

solve x - y ln y dx + y ln x dx + x ln y dy - x ln x dy = 0

OpenStudy (isabel☺):

i was stuck :( i am trying to integrate: \[\int\limits \ln (vx) dv\]

OpenStudy (dls):

attempt?

OpenStudy (goformit100):

Show your prior working.

hartnn (hartnn):

\(\int \ln (vx) dv\) here, x is a constant, so its just like \(\int \ln (ax)dx\)

hartnn (hartnn):

and ofcourse, you will need uv rule

OpenStudy (anonymous):

\[ \begin{split} \int \ln(vx)dv &= \int \ln(vx)\frac {xdv}x \\ &= \int \ln(vx)\frac {d(xv)}x \end{split} \]At this point it is pretty simple.

hartnn (hartnn):

x dv = d(xv) ?? :O i though d(xv) = x dv+ v dx

OpenStudy (anonymous):

Well, I treated \(x\) like a constant...

OpenStudy (anonymous):

So that means \(v~dx = 0\).

hartnn (hartnn):

oh...you still need uv rule

OpenStudy (anonymous):

You can use product rule on a constant, it's just that the derivative of a constant is 0 so only one term actually remains.

OpenStudy (anonymous):

\(dx =0\).

OpenStudy (isabel☺):

i don't get the part where xdv=d(vx)

OpenStudy (anonymous):

Well: \[ d(xy) = \frac{d(xv)}{dv} dv \]This is just a definition thing. Anyway, since \(d(xv)/dv = x\) we get: \[ d(xy) = x~dv \]

OpenStudy (anonymous):

In general: \[ dy = \frac {dy}{dx}dx \]

OpenStudy (anonymous):

You are probability familiar with:\[ \frac{dy}{dx} = \lim_{\Delta x\to 0}\frac{\Delta y}{\Delta x} = \lim_{\Delta x\to 0}\frac{y(x+\Delta x)-y(x)}{\Delta x} \]Basically, \(d\) in this case is notation convenience for \(\Delta \) under some limit.

OpenStudy (anonymous):

To avoid having to write out the variable of differentiation, we can define differentials as: \[ dy = \frac{dy}{dx}dx \]and \(x\) can be any variable in this case. When we use substitution, or chain rule, we are just changing the variable of differentiate: \[ \int f(g(x))g'(x)~dx = \int f(g)~dg, \quad dg = g'(x)dx \]

OpenStudy (isabel☺):

simpler solution please? sorry i still don't get that :(

hartnn (hartnn):

you know uv rule of integration ? can you solve your integral with that rule ?

OpenStudy (isabel☺):